OpenStudy (anonymous):

limit 3x/5sin10x x-->o

OpenStudy (anonymous):

ok since 3x tends to 0 as x tends to zero, and also 5sin(10x) tends to zero as x tends to zero, and they are both differentiable functions, we may apply L'Hopital's Rule

OpenStudy (anonymous):

\[\lim_{x\rightarrow 0} \frac{3x}{5\sin (10x)} = \lim_{x\rightarrow 0} \frac{3}{50\cos (10x)} = \frac{3}{50}\]